OpenStudy (anonymous):

Shannon is managing a small zoo and has been analyzing the attendance data. Shannon has found that the number of sleeping animals increases exponentially as the temperature increases. Shannon has also found a linear equation that models the number of people who leave the park early depending on the number of sleeping animals. Describe how Shannon can combine these two functions into a new function and explain what that function would predict. @wolf1728

OpenStudy (wolf1728):

Geez that gives you no numbers to work with.

OpenStudy (epicteatime):

:P

OpenStudy (anonymous):

yeah i know

OpenStudy (wolf1728):

Well we can think up something. Above 80° 1 animal asleep Above 85° 2 animals asleep Above 90° 4 animals asleep Above 95° 8 animals asleep Hmm there's a pattern developing. sleeping animals = 2^ (Air Temp - Temp 80° F) Air temp 80° - then 2^(80-80) = 2^0 = 1

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@wolf1728 what happened

OpenStudy (wolf1728):

For the linear equation, we could say for every sleeping animal, 2 people leave the park early. 2 animals asleep, 4 people leave, 3 animals asleep 6 people leave, etc. I'm working on the linear equation.

OpenStudy (wolf1728):

People leaving the park early = 2 * ((Air Temp -75°) / 5)

OpenStudy (anonymous):

So thats it?

OpenStudy (wolf1728):

No - the remaining thing to do is : "Describe how Shannon can combine these two functions into a new function and explain what that function would predict."

OpenStudy (anonymous):

Thank You

OpenStudy (wolf1728):

Do you think you could do that?

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

Good - well I think I will leave Open Study for a while - see you woohoo.